OpenStudy (anonymous):

i^154=? explanation please

OpenStudy (anonymous):

\[i^{154}=(i^4)^{38}\times i^2=1^{38}\times-1=-1\]

OpenStudy (anonymous):

ummm.....

OpenStudy (amistre64):

there are only 4 i's , everything else is a modular of them

OpenStudy (anonymous):

modular?

OpenStudy (amistre64):

154/4 = 38 with a remainder of 2 i^2 = -1

OpenStudy (anonymous):

amistre64 can you please check my problems to make sure that I got the correct answer please. Sorry to interrupt...

OpenStudy (anonymous):

thanks amister64 would you be able to answer any other question i'd have?

OpenStudy (amistre64):

modular is just clock addition; 9oclock + 6 hours is 15oclock; but its says 3 clock why? 15/12 = 1 remainder 3 ..... 3 oclock

OpenStudy (amistre64):

i can try

OpenStudy (anonymous):

ok thanks